OpenStudy (anonymous):

the binomials (x-2) and (2x+3) are the factors of which polynomial ? A) 2x^2-6 ; B) 2x^2-x-6 ; C)2x^2+x-6 OR D) 2x^2+7x-6 ?

OpenStudy (raden):

just multiply of (x-2)(2x+3), can you expand this ?

OpenStudy (anonymous):

idk the choices are A) 2x^2-6 B) 2x^2-x-6 C) 2x^2+x-6 D) 2x^2+x-6

OpenStudy (raden):

use the FOIL method (x-2)(2x+3) multiplying the First terms, x * 2x = 2x^2 multiplying the Outside terms, x * 3 = 3x multiplying the Inside terms, -2 * 2x = -4x, and multiplying the Last terms, -2 * 3 = -6 ow add them, becomes 2x^2 + 3x + (-4x) + (-6) = 2x^2 -x -6

OpenStudy (raden):

* now

OpenStudy (anonymous):

ohhh I think its 2x^2-x-6

OpenStudy (raden):

yes

OpenStudy (raden):

this site might useful http://www.mathwords.com/f/foil_method.htm

OpenStudy (anonymous):

okay thank youu

OpenStudy (raden):

you're welcome

OpenStudy (anonymous):

I need your help again

OpenStudy (raden):

i'll try, what your question

OpenStudy (anonymous):

which is NOT a factor of 2x^2-32 ? A) x^2-32 B) 2 C) x-4 D) x^2-16

OpenStudy (raden):

the factors of 2x^2-32 are 2(x^2 - 16) = 2(x-4)(x+4) so,we have the factors : 2, x-4, x+4, x^2-16 i didnt see the x^2-32 be one of the factor